OpenStudy (anonymous):

What is a possible first step to eliminate the variable x in the following system? 2x - 7y = 3 4x + 5y = -2

OpenStudy (anonymous):

you want to get 2x and 4 x to cancel right so one has to be positive and one negative right?

OpenStudy (anonymous):

\[-2(2x-7y=3)\] \[4x+5y=-2\]

OpenStudy (anonymous):

now if you distribute the -2 to the top equation what do you get?

OpenStudy (anonymous):

@danny562 does this make sense ?

OpenStudy (anonymous):

a) Multiply the first equation by -3 and the second equation by 2 b) Multiply the first equation by 3 and the second equation by -2 c) Multiply the first equation by 4 and the second equation by 3 d) All of the above

OpenStudy (anonymous):

\[-4x+14y=-6\] \[4x+5y=-2\] none of those make the x's cancel are you sure the problem is copied right ?